A cup is made from 200 g of silver and contains 200 g of water. Both cup and water are at room temperature (24oC). A 40-g piece of unknown material is heated to 85oC and then added to the water. The heat released to the environment by the cup, water, and object system is negligible. After achieving thermal equilibrium, all three objects have a temperature of 25oC. What is the specific heat of the unknown material?1.      For water, c = 4186 J/kg.Co. For silver, c = 235 J/kg.Co

1
Expert's answer
2022-01-26T07:40:44-0500
"c_sm_s(t_f-t_r)+c_wm_w(t_f-t_r)+cm(t_f-t)=0,\\\space\\
c=\dfrac{c_sm_s(t_f-t_r)+c_wm_w(t_f-t_r)}{m(t-t_f)},\\
c=368\text{ J}/\text{(kg}·\text K)."